### v1

@@ 1 Introduction
Large language models (LLMs) are revolutionizing research in various X+Language tasks (with X be vision, robotics, sciences, etc.), ushering in a new era with unprecedented advancements in computer science (OpenAI, 2023; Chung et al., 2022; Touvron et al., 2023; Chiang et al., 2023; Dubey et al., 2024). Video Question Answering (VideoQA), as a multimodal and multifaceted challenge that requires both linguistic and visual reasoning, is at the forefront of this revolution. It also serves as a litmus test for evaluating the comprehensive multimodal understanding and reasoning capabilities of cutting-edge Video-LLMs 111In this paper, we denote Video-LLMs as video-language models that use LLMs with billion parameters. (Alayrac et al., 2022; Dai et al., 2023; Maaz et al., 2023; Chen et al., 2023; Liu et al., 2023; Zhang et al., 2023c, b; Lin et al., 2023; Li et al., 2023e; Team et al., 2023).
@@
We begin by collecting a series of representative Video-LLMs (Yang et al., 2022; Ko et al., 2023; Yu et al., 2023; Zhang et al., 2023a, b; Li et al., 2023e; Lin et al., 2023; Li et al., 2024a). These models have achieved state-of-the-art (SOTA) on VideoQA in either finetuned or zero-shot settings by harnessing popular LLMs such as DeBERTa-V2 (He et al., 2020), Flan T5 (Chung et al., 2022), LLaMA (Touvron et al., 2023), Vicuna (Chiang et al., 2023) and GPT-4 (OpenAI, 2023). The accuracy of these models reaches around 70%80% on popular benchmarks such as NExT-QA (Xiao et al., 2021) and MSVD-QA (Xu et al., 2017), for both multi-choice (MC) QA and open-ended (OE) QA.
@@ 2.2 VideoQA Technique Evolution
The VideoQA techniques generally follow the corresponding development of vision and text representation models. The development can be roughly divided into three stages. The first is 1) CNN + RNN (Jang et al., 2017; Xu et al., 2017; Le et al., 2020; Li et al., 2022; Xiao et al., 2022a) where video and text are encoded by convolutional neural networks (e.g., ResNet (He et al., 2016)) and recurrent neural networks (e.g., LSTM (Hochreiter and Schmidhuber, 1997)) respectively. The second stage 2) CNN/ViT+BERT (Sun et al., 2019; Li et al., 2020; Seo et al., 2021; Lei et al., 2021; Yang et al., 2021; Xiao et al., 2022b, 2023; Fu et al., 2023; Li et al., 2023g, f), features improved visual (Dosovitskiy et al., 2021) and language representations (Devlin et al., 2018; Liu et al., 2019), where the models also benefit from self-supervised cross-modal pre-training at scale and fine-tuned small language models (e.g., BERT and RoBERTa) on the target datasets. Finally, the third (and the current) stage 3) CLIP+LLM (Yang et al., 2022; Liu et al., 2023; Maaz et al., 2023; Li et al., 2023a; Yu et al., 2023; Ko et al., 2023; Zhang et al., 2023b; Dai et al., 2023; Zhang et al., 2023c; Lin et al., 2023) features pre-trained cross-modal visual encoders (e.g., CLIP-ViT (Radford et al., 2021)) and frozen LLMs (e.g., Flan T5 (Chung et al., 2022), LLaMA (Touvron et al., 2023)) with instruction tuning of projection and adaptation modules. In the transition of each stage, there are significant performance jumps. Especially in the LLM stage, leading models like GPT-4V/o (OpenAI, 2023) and Gemini (Team et al., 2023) rival human performance in answering standard visual questions.
@@ 2.3 LLMs for VideoQA
Existing LLM-based VideoQA methods can be roughly classified into 3 groups: 1) General-purpose MLLMs, where models are not specialized to VideoQA and are instead capable of handling various vision-language tasks such as image/video QA and captioning. Typical models are InstructBLIP (Dai et al., 2023), Video-ChatGPT (Maaz et al., 2023), Video-LLaMA (Zhang et al., 2023b), VideoChat (Li et al., 2023b), VideoChat2 (Li et al., 2024a), Video-LLaVA (Lin et al., 2023), LLaMA-VID (Li et al., 2023e), and LLaVA-NeXT (Zhang et al., 2024e), and commercial models such as GPT-4V/4o (OpenAI, 2023), Gemini (Team et al., 2023). These models do not finetune on the target datasets but are instruction tuned and prompted to perform different tasks, e.g. VideoQA. We refer to the recent survey (Tang et al., 2023) for a more detailed introduction.
@@
3) Tool-based Video-LLMs. These methods are parameter-free and focused on the use or collaboration of different MLLMs (LLaVA, GPT-4V, LaViLA (Zhao et al., 2023)) and LLMs (LLaMA, GPT-3.5 and GPT-4) as tools (e.g., planning, captioning, and question-answering) for VideoQA. Representative examples include ViperGPT (Surís et al., 2023), MoReVQA (Min et al., 2024), LLoVi (Zhang et al., 2023a), IG-VLM (Kim et al., 2024), TraveLER (Shang et al., 2024), VideoAgent (Wang et al., 2024a; Fan et al., 2024) and VideoTree (Wang et al., 2024b). These methods are training-free and perform zero-shot VideoQA.
@@ 2.4 Empirical VQA Study
| LLaMA-VID (Li et al., 2023e) | - | CLIP (Radford et al., 2021) | Vicuna (7B, 13B) (Chiang et al., 2023) | |
@@ 3.4 Temporal Understanding
Among the models, GPT-4o significantly outperforms others in the TD test even though it does not encode any temporal information (we specify in the prompts that the images are ordered video frames). By jointly considering the accuracy in Fig. LABEL:fig:temp-acc-res, we speculate that the models trained on the target datasets (all models except for GPT-4o and LLoVi (Zhang et al., 2023a)) likely overfit to dataset statistics rather than reason about time faithfully. Additionally, LLaMA-VQA (Ko et al., 2023) shows the smallest flip rate in TE test. We attribute such strength of LLaMA-VQA to its auxiliary training objective by inversely predicting the questions given the answers, which implicitly suits the TE probes by enhancing the relation between the answer and the contextual term in the question. Moreover, Fig. LABEL:fig:temp-acc-res shows that models using advanced LLMs such as LLaMA (Li et al., 2023e) and Vicuna (Chiang et al., 2023) perform better on the TE test but worse on the TD test than non-LLM method CoVGT. This indicates that LLM-based methods are better at exploiting language priors or commonsense in the TE questions for predicting answers, even though they seem weaker than non-LLM methods in understanding video content order. Again, we find that GPT-4o achieves the highest accuracy in both TE and TD tests, showcasing strong zero-shot temporal QA capability.
@@ 3.5 Visual Grounding
2) Compared with the non-LLM method (CoVGT), specilized Video-LLMs (e.g., FrozenGQA and SeViLA) benefit more from short-cut learning. This is because they remarkably surpass the non-LLM method on normal VQA accuracy (by 10%) but are only marginally better on GQA (e.g., by 2%) (Fig. LABEL:fig:ground_nbp). Also their LLMs alone (e.g., DeBERTa-V2-XL and Flan-T5-XL) already wins over CoVGT’s language model (RoBERTa) by more than 7% according to the corresponding BlindQA accuracies. A further comparison of NormalVQA vs. BlindQA (Fig. LABEL:fig:ground_nbp), NormalVQA vs. PosVQA vs. NegVQA (Fig. LABEL:fig:ground_nnp) indicates that the short-cuts are primarily from language priors and spurious vision-text correlations. For example, the BlindQA baselines without any visual inputs retain more than 70% of the corresponding NormalVQA’ performance. Moreover, visual signals that come from answer-irrelevant video segments yield surprisingly high performance that is competitive to model variants with answer-relevant (ground-truth) visual segments (NegVQA vs. PosVQA).
@@
Additionally, in our implementation, when GPT-4o is prompted to also output the indexes of the video frames that it uses to predict the answers versus just answering the questions. Its QA accuracy increases remarkably by more than 15%, from 49.0% to 64.7% for NormalVQA. This suggests GPT-4o’s great potential in reasoning like human and providing visually grounded answers if appropriately prompted. We believe this is also an interesting finding and raise it (how to use/prompt MLLMs like GPT-4o) as an open challenge for future exploration.
@@ 3.8.2 Generalization to OOD data
4 Discussion
We summarize the major limitations of Video-LLMs and discuss possible directions for improvement.
Robustness and Trustworthiness. All probes showed that the open-sourced Video-LLMs be it general-purpose or fine-tuned, perform well for standard question-answering, but suffer under adversarial language modifications while being unaffected by analogous video modifications. These findings reveal a lack of robustness and interpretability and underscores the need for rationales in Video-LLM developing. Some recent studies in multimodal chain-of-thought (Wei et al., 2022; Zhang et al., 2023e; Fei et al., 2024) and modular multi-agent approaches (Min et al., 2024; Shang et al., 2024) show improved robustness and interpretability. Yet, one needs to mind the inefficiency issues of such techniques. In addition, video multimodal benchmarks are emerging (Li et al., 2024a; Fu et al., 2024; Liu et al., 2024b; Li et al., 2024b), but almost all of them target comprehensive video understanding, with little attention on interpretability. Related benchmarks such as VIP (Himakunthala et al., 2023) and NExT-GQA (Xiao et al., 2024) are thus valuable efforts in developing trustworthy Video-LLM techniques, and deserve more research efforts.
@@ 4 Discussion
Efficient and Fine-Grained Long-Form Video Understanding. Our observations about GPT-4o’s behavior in VideoQA (refer to Fig. 1 and 5), as well as the rapid evolution of SOTA results on common VideoQA benchmarks (e.g., MSVD-QA) suggests that coarse and holistic understanding of short videos (less than 20s) are solvable by scaling up existing Video-LLMs or operating powerful image MLLMs (like GPT-4o) frame by frame. Fine-Grained Long-Form Video Understanding manifest two major challenges for existing Video-LLMs:
@@
The first is a ‘needle in the haystack’ challenge - search in the long video for short moments (or key frames) that carry the correct answer. Existing benchmarks such as NExT-GQA (Xiao et al., 2024) and Ego4D-QA (Datta et al., 2022) embody such challenges, yet effective techniques are lacking. Recent solutions (e.g., MoReVQA (Min et al., 2024), VideoAgent (Wang et al., 2024a; Fan et al., 2024) and TraveLER (Shang et al., 2024)) show promise, but they are inefficient, brute-force methods that traverse the whole video frame-by-frame (or segment-by-segment) and perform LLM inference multiple times. Thus, efficiently understanding long videos is still an open problem.
@@
The second challenge is fine-grained understanding of very long-form videos, with large temporal certificates, as defined in EgoSchema (Mangalam et al., 2023). Our studies in Sec. 3.8.2 show that existing Video-LLMs targeting short clips do not generalize well to such videos. The aforementioned traversal solutions are ineffective as the task requires reasoning over and weaving together multiple frames or segments to achieve holistic understanding. Recent memory- and caption-based solutions (e.g., MoReVQA (Min et al., 2024), LLoVi (Zhang et al., 2023a) and VideoAgent (Fan et al., 2024)) as well as the direct long-frame-encoding solutions (e.g., LongVA (Zhang et al., 2024a)) offer some insights to handle long videos, but understanding the videos in efficient and fine-grained manner remains a key challenge.
@@
Egocentric and Embodied QA. We primarily analyzed VideoQA with datasets featuring third-person view (exo-centric) videos and QAs, be it effective as benchmark evaluation for Video-LLMs. Yet noteworthy an important real-world application for VideoQA is geared towards reliable embodied QA assistance (Grauman et al., 2022; Majumdar et al., 2024). In this regard, first-person view (ego-centric) VideoQA are of greater interest and relevance. Sec. 3.8.2 shows that Video-LLMs pre-trained on third-person view VideoQA datasets can generalize to these application scenarios, but the performance are weak compared with the tool-based Video-LLMs (Zhang et al., 2023a) that directly use ego-centric pretrained visual descriptions (Zhao et al., 2023), as well as the well-prompted general-purpose MLLMs (e.g. GPT-4o) (Liu et al., 2024a). Therefore, the efforts of leveraging existing MLLMs for ego-centric VideoQA will be of critical importance to initiate successful embodied assistance applications.

### v2

@@ 1 Introduction
Large language models (LLMs) are revolutionizing research in various X+Language tasks (with X be vision, robotics, sciences, etc.), ushering in a new era with unprecedented advancements in computer science (OpenAI, 2023; Chung et al., 2022; Touvron et al., 2023; Chiang et al., 2023; Dubey et al., 2024). Video Question Answering (VideoQA), as a multimodal and multifaceted challenge that requires both linguistic and visual reasoning, is at the forefront of this revolution. It also serves as a litmus test for evaluating the comprehensive multimodal understanding and reasoning capabilities of cutting-edge Video-LLMs 111In this paper, we denote Video-LLMs as video-language models that use LLMs with billion parameters. (Alayrac et al., 2022; Dai et al., 2023; Maaz et al., 2023; Chen et al., 2023; Liu et al., 2023; Zhang et al., 2023b; Lin et al., 2023; Li et al., 2024c; Team et al., 2023).
@@
We begin by collecting a series of representative Video-LLMs (Yang et al., 2022; Ko et al., 2023; Yu et al., 2023; Zhang et al., 2023a, b; Li et al., 2024c; Lin et al., 2023; Li et al., 2024a). These models have achieved state-of-the-art (SOTA) on VideoQA in either finetuned or zero-shot settings by harnessing popular LLMs such as DeBERTa-V2 (He et al., 2020), Flan T5 (Chung et al., 2022), LLaMA (Touvron et al., 2023), Vicuna (Chiang et al., 2023) and GPT-4 (OpenAI, 2023). The accuracy of these models reaches around 70%80% on popular benchmarks such as NExT-QA (Xiao et al., 2021) and MSVD-QA (Xu et al., 2017), for both multi-choice (MC) QA and open-ended (OE) QA.
@@ 2.2 VideoQA Technique Evolution
The VideoQA techniques generally follow the corresponding development of vision and text representation models. The development can be roughly divided into three stages. The first is 1) CNN + RNN (Jang et al., 2017; Xu et al., 2017; Le et al., 2020; Li et al., 2022; Xiao et al., 2022a) where video and text are encoded by convolutional neural networks (e.g., ResNet (He et al., 2016)) and recurrent neural networks (e.g., LSTM (Hochreiter and Schmidhuber, 1997)) respectively. The second stage 2) CNN/ViT+BERT (Sun et al., 2019; Li et al., 2020; Seo et al., 2021; Lei et al., 2021; Yang et al., 2021; Xiao et al., 2022b, 2023; Fu et al., 2023; Li et al., 2023f, e), features improved visual (Dosovitskiy et al., 2021) and language representations (Devlin et al., 2018; Liu et al., 2019), where the models also benefit from self-supervised cross-modal pre-training at scale and fine-tuned small language models (e.g., BERT and RoBERTa) on the target datasets. Finally, the third (and the current) stage 3) CLIP+LLM (Yang et al., 2022; Liu et al., 2023; Maaz et al., 2023; Li et al., 2023a; Yu et al., 2023; Ko et al., 2023; Zhang et al., 2023b; Dai et al., 2023; Zhang et al., 2023c; Lin et al., 2023) features pre-trained cross-modal visual encoders (e.g., CLIP-ViT (Radford et al., 2021)) and frozen LLMs (e.g., Flan T5 (Chung et al., 2022), LLaMA (Touvron et al., 2023)) with instruction tuning of projection and adaptation modules. In the transition of each stage, there are significant performance jumps. Especially in the LLM stage, leading models like GPT-4V/o (OpenAI, 2023) and Gemini (Team et al., 2023) rival human performance in answering standard visual questions.
@@ 2.3 LLMs for VideoQA
Existing LLM-based VideoQA methods can be roughly classified into 3 groups: 1) General-purpose MLLMs, where models are not specialized to VideoQA and are instead capable of handling various vision-language tasks such as image/video QA and captioning. Typical models are InstructBLIP (Dai et al., 2023), Video-ChatGPT (Maaz et al., 2023), Video-LLaMA (Zhang et al., 2023b), VideoChat (Li et al., 2023b), VideoChat2 (Li et al., 2024a), Video-LLaVA (Lin et al., 2023), LLaMA-VID (Li et al., 2024c), and LLaVA-NeXT (Zhang et al., 2024e), and commercial models such as GPT-4V/4o (OpenAI, 2023), Gemini (Team et al., 2023). These models do not finetune on the target datasets but are instruction tuned and prompted to perform different tasks, e.g. VideoQA. We refer to the recent survey (Tang et al., 2023) for a more detailed introduction.
@@
3) Tool-based Video-LLMs. These methods are parameter-free and focused on the use or collaboration of different MLLMs (LLaVA, GPT-4V, LaViLA (Zhao et al., 2023)) and LLMs (LLaMA, GPT-3.5 and GPT-4) as tools (e.g., planning, captioning, and question-answering) for VideoQA. Representative examples include Socratic Models (Zeng et al., 2023), ViperGPT (Surís et al., 2023), MoReVQA (Min et al., 2024), LLoVi (Zhang et al., 2023a), IG-VLM (Kim et al., 2024), TraveLER (Shang et al., 2024), VideoAgent (Wang et al., 2024a; Fan et al., 2024) and VideoTree (Wang et al., 2024b). These methods are training-free and perform zero-shot VideoQA.
@@ 2.4 Empirical VQA Study
| LLaMA-VID (Li et al., 2024c) | - | CLIP (Radford et al., 2021) | Vicuna (7B, 13B) (Chiang et al., 2023) | |
@@ 3.4 Temporal Understanding
Among the models, GPT-4o significantly outperforms others in the TD test even though it does not encode any temporal information (we specify in the prompts that the images are ordered video frames). By jointly considering the accuracy in Fig. LABEL:fig:temp-acc-res, we speculate that the models trained on the target datasets (all models except for GPT-4o and LLoVi (Zhang et al., 2023a)) likely overfit to dataset statistics rather than reason about time faithfully. Additionally, LLaMA-VQA (Ko et al., 2023) shows the smallest flip rate in TE test. We attribute such strength of LLaMA-VQA to its auxiliary training objective by inversely predicting the questions given the answers, which implicitly suits the TE probes by enhancing the relation between the answer and the contextual term in the question. Moreover, Fig. LABEL:fig:temp-acc-res shows that models using advanced LLMs such as LLaMA (Li et al., 2024c) and Vicuna (Chiang et al., 2023) perform better on the TE test but worse on the TD test than non-LLM method CoVGT. This indicates that LLM-based methods are better at exploiting language priors or commonsense in the TE questions for predicting answers, even though they seem weaker than non-LLM methods in understanding video content order. Again, we find that GPT-4o achieves the highest accuracy in both TE and TD tests, showcasing strong zero-shot temporal QA capability.
@@ 3.5 Visual Grounding
2) Compared with the non-LLM method (CoVGT), specialized Video-LLMs (e.g., FrozenGQA and SeViLA) benefit more from short-cut learning. This is because they remarkably surpass the non-LLM method on normal VQA accuracy (by 10%) but are only marginally better on GQA (e.g., by 2%) (Fig. LABEL:fig:ground_nbp). Also their LLMs alone (e.g., DeBERTa-V2-XL and Flan-T5-XL) already wins over CoVGT’s language model (RoBERTa) by more than 7% according to the corresponding BlindQA accuracies. A further comparison of NormalVQA vs. BlindQA (Fig. LABEL:fig:ground_nbp), NormalVQA vs. PosVQA vs. NegVQA (Fig. LABEL:fig:ground_nnp) indicates that the short-cuts are primarily from language priors and spurious vision-text correlations. For example, the BlindQA baselines without any visual inputs retain more than 70% of the corresponding NormalVQA’ performance. Moreover, visual signals that come from answer-irrelevant video segments yield surprisingly high performance that is competitive to model variants with answer-relevant (ground-truth) visual segments (NegVQA vs. PosVQA).
@@
Additionally, in our implementation, when GPT-4o is prompted to also output the indexes of the video frames that it uses to predict the answers versus just answering the questions, its QA accuracy increases remarkably by more than 15%, from 49.0% to 64.7% for NormalVQA. This suggests GPT-4o’s great potential in reasoning like human and providing visually grounded answers if appropriately prompted. We believe this is also an interesting finding and raise it (how to use/prompt MLLMs like GPT-4o) as an open challenge for future exploration.
@@ 3.8.2 Generalization to OOD data
4 Discussion & Future Directions
4.1 Robustness and Trustworthiness
All our probes showed that the open-sourced Video-LLMs perform well for standard question-answering. However, these models are susceptible to adversarial language modifications while remain stubbornly unaffected to analogous video modifications. These findings reveal a lack of robustness, underscoring the urgent need for interpretability in Video-LLM development. There are many directions for improvement. We outline some of them, dividing them up from the perspective of data, modeling, and learning.
From a data perspective, a direct solution could be to apply our strategies of designing adversarial probes to augment the training data. Our simple attempt (see Appendix .5) shows that this can effectively improve the model performance on the probe data and the original test data. Aside from training data, new video testing benchmarks are emerging (Li et al., 2024a; Fu et al., 2024; Liu et al., 2024b; Li et al., 2024b). Yet almost all of them target comprehensive video understanding, with little regard for interpretability. Related benchmarks such as VIP (Himakunthala et al., 2023) and NExT-GQA (Xiao et al., 2024) are thus valuable efforts in developing trustworthy Video-LLM techniques and deserve more research attention.
From a modeling perspective, VideoQA models are imbalanced in terms of their language and visual representations. The language encoders are drawn from ever-larger and more powerful language models which dominate over comparatively weaker visual representations, where CLIP Radford et al. (2021) and its variants Ilharco et al. (2021); Sun et al. (2023) remain de-facto standards even though they struggle to represent fine-grained details. More importantly, from an image to a video level, the temporal and dynamic modeling tends to be primitive. Hence, exploring and improving the image or video encoders would contribute to more robust models.
Finally, from a learning perspective, existing models learn to directly output the final answers in an auto-regressive manner. Directly targeting the answers (especially lengthy ones) enables the models to ignore the visual contents and take whatever shortcuts they can (e.g., hallucinate based on commonsense knowledge from LLMs) to make plausible but often incorrect responses (Sec. 3.7). As such, we believe the focus on intermediate reasoning steps and post-verifications can benefit model developing towards robustness and trustworthiness. Some recent studies in multimodal chain-of-thought (Wei et al., 2022; Zhang et al., 2023e; Fei et al., 2024) and modular multi-agents (Min et al., 2024; Shang et al., 2024) provide insights in this direction and show improved robustness and interpretability.
4.2 Efficient and Fine-Grained Long-Range Understanding
@@
GPT-4o’s superior behavior in VideoQA (see Fig. 1 and 5), as well as the ascent of SOTA performance on benchmarks like MSVD-QA and MSRVTT-QA suggests that coarse and holistic understanding of short videos (less than 20s) are solvable by scaling up existing Video-LLMs or prompting powerful image MLLMs (like GPT-4o). Yet the understanding of fine-grained details (Xiao et al., 2021) or long-range videos (Mangalam et al., 2023; Fu et al., 2024) remains out of reach. These videos offer additional challenges that require more dedicated analysis and architecture design.
@@
One such hurdle comes from the ‘needle in the haystack’ problem - the search in the long video for short moments (or keyframes) that carry the correct answer. Existing benchmarks such as NExT-GQA (Xiao et al., 2024) and QAEgo4D (Datta et al., 2022) embody such challenges, yet effective techniques are lacking. Recent solutions (e.g., MoReVQA (Min et al., 2024), VideoAgent (Wang et al., 2024a; Fan et al., 2024) and TraveLER (Shang et al., 2024)) show promise. However, they are relatively inefficient, in that they traverse the entire video frame-by-frame, or segment-by-segment, and perform LLM inference multiple times. Efficient understanding of such long videos is gaining attention but continues to be an open problem for the foreseeable future.
@@
The second challenge is a fine-grained understanding of long-range videos. The concept of large temporal certificates, as defined in EgoSchema (Mangalam et al., 2023), highlights accurately differentiating the visually very similar long activities (e.g., teaching someone to play vs. play with someone). Our studies in Sec. 3.8.2 show that existing Video-LLMs targeting short clips do not generalize well to such videos. Traversal-based solutions are ineffective as the task requires reasoning over and weaving together (potentially distant) frames or segments to achieve holistic understanding. Recent memory- and caption-based solutions (e.g., MoReVQA (Min et al., 2024), LLoVi (Zhang et al., 2023a) and VideoAgent (Fan et al., 2024)) as well as the direct long-range encoding solutions (e.g., LongVA (Zhang et al., 2024a)) offer some insights. However, fine-grained understanding, especially in an efficient manner, remains a key challenge.
4.3 Egocentric and Embodied QA
@@
We primarily analyzed VideoQA with exo-centric datasets - videos from the third-person view. It should be noted that an important real-world application for VideoQA is geared towards reliable embodied assistance (Grauman et al., 2022; Majumdar et al., 2024). In this regard, ego-centric, or first-person view videos and associated question-answer pairs are of great interest and relevance. Sec. 3.8.2 shows that Video-LLMs pre-trained on third-person view VideoQA datasets can generalize to these application scenarios. However, their performance is weak compared with the tool-based Video-LLMs (Zhang et al., 2023a) that directly use ego-centric pretrained visual descriptions (Zhao et al., 2023), and well-prompted general-purpose MLLMs (e.g. GPT-4o) (Liu et al., 2024a). Therefore, the efforts of leveraging existing MLLMs for ego-centric VideoQA will be of critical importance in initiating successful embodied assistance applications.